Discrimination. Let Wf denote wages actually paid to women, Wm
wages actually paid to men, and d the discrimination coefficient. Assume
the type of (employer) discrimination analyzed by Becker where firms act
as is they pay Wf (1 + d) when they hire women and Wf when they hire
men.
(a) Define discrimination in the labour market.
(b) In the case where the value of the marginal products of male and
female labour are the same, solve for the ratio of female to male
wages when the discrimination coefficient is 0.20.
(c) What is the wage gap, due to discrimination, in this case?
(d) Explain Becker’s theory. What is meant by a discrimination coefficient
that is 0.20?
1
Expert's answer
2017-04-07T10:34:07-0400
Firms pay Wf (1 + d) when they hire women and Wf when they hire men.
(a) Discrimination in the labour market occurs, when men have more employment possibilities and are paid more than women.
(b) In the case where MPm = MPf, then ratio of female to male wages when d = 0.20 is:
MPm/Wm = MPf/(Wf*1.2),
MPm/MPf = Wm/(Wf*1.2),
Wm/Wf = 1.2.
(c) The wage gap, due to discrimination, in this case is 20%.
(d) Becker recognized that people (employers, customers, and employees) sometimes do not want to work with minorities because they have preference against the disadvantaged groups.